OpenStudy (anonymous):

If an object is dropped from a height of 144 feet, the function h(t)=16t^2+144 gives the height of the object after t seconds. When will the object hit the ground? A. 9 s B. 6 s C. 1.5 s D. 3 s

OpenStudy (anonymous):

(B)

OpenStudy (anonymous):

presumably it is \(h(t)=-16t^2+144\) set \(-16t^2+144=0\) solve in 3 steps

OpenStudy (anonymous):

\[16t^2=144\] divide by 16, take the square root

OpenStudy (anonymous):

9

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[144\div 16=9\] so you have \(t^2=9\) but you want \(t\) not \(t^2\) what number squared is 9 ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

you win

OpenStudy (anonymous):

(D)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw